solve for y when x =2
2y(4-x)=x/2

Answer by algebrahouse.com(1659)

2y(4 - x) = x/2.....when x = 2
2y(4 - 2) = 2/2 {substituted 2 in for x}
2y(2) = 1 {simplified}
4y = 1 {multiplied 2y by 2}
y = 1/4 {divided both sides by 4}

For more help from me, visit: www.algebrahouse.com